OpenStudy (anonymous):

Find the midpoint of segment TS with endpoints T(-6, 2) and S(3, 4). A. (-4.5, 3) B. (-1.5,3) C. (-1, -1.5) D. (4.5, -3)

OpenStudy (aaliyahfairgood):

TS= square root of (-6-3)squared +(2-4)squared TS=square root of (-9)squared +(-2)squared TS=square root of (81+4) TS= Square root of 85

OpenStudy (anonymous):

Well, you're going to have to use the midpoint formula. Do you know what that is? o.e

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@vangoo98

OpenStudy (aaliyahfairgood):

yeah just a sec

OpenStudy (anonymous):

K, well the midpoint formula is:\[M=(x1+x2/2, y1+y2/2)\]

OpenStudy (aaliyahfairgood):

Sorry I gave you distance first Midpoint = (x1+x2/2) , (y1+y2/2) So since TS is T(-6, 2) and S(3, 4). TS= (-6+3/2), (2+4/2) TS= (-3/2),(6/2) TS= (-3/2),(3)

OpenStudy (anonymous):

So, given the two points, -6 will be your x1 and 3 will be your x2. Similarly, 2 will be your y1 and 4 will be your y2. Now, just plug them into the equation: It would be (-6+3/2, 2+4/2) (-1.5,3) is your answer. So, B. Hope this helped :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem! :)